OpenStudy (anonymous):

For the Experiment, flip a coin until heads shows, assume that the probability of heads on one flip is 3/5 we define a RV (random variable) X= the number of flip. 1)what are the possible values of X. Find the probability distribution function for X: gives the first four values, and then find a general formula for the probability that x=n.

OpenStudy (anonymous):

you could get it on the first try, the second try, the third try etc \[X=1,2,3,...\]

OpenStudy (anonymous):

p(x=1)= 3/5 p(x=2)= 2/5*3/5 p(x=3)= (2/5)^2* 3/5 ?

OpenStudy (anonymous):

@satellite73 that all i need to do ?